A 0.1375-gram sample of solid magnesium is burned in a constant-volume bomb calorimeter that has a heat capacity of 3024 J/⁰C . The temperature increases by 1.126 ⁰C. Calculate the heat given off by the burning of Mg, in kJ/g and in kJ/mol. (Atomic Mass of Mg = 24.31 g/mol)


1
Expert's answer
2020-09-29T06:53:20-0400

"Q" cal "= C"cal"\Delta T"

"=3024 \times 1.126"

"=3405J"

then heat of reaction of magnesium solid in "kj\/j"

"Q"rn "= \frac{(3405\times .1375)}{1000}"

"= -24.76kj\/g"

so that amount of heat liberated in reaction is "= -24.76kj\/g"


"Q"rn "= -\frac{24.76kj}{gMg} \times \frac{24.31gMg}{1MOL Mg}"

"= - 601.9 kj\/mol"


so that amount of heat liberated in reaction is"= - 601.9 kj\/mol"